OpenStudy (anonymous):

A rectangular garden is 5 ft longer than it is wide. Its area is 336 ft2. What are its dimensions? --- find its length n width

OpenStudy (anonymous):

i would use substitution method

OpenStudy (wolf1728):

l * w = 336 w + 5 = L W = L-5 W= 336 /L L-5 = 336/L L² - 5L = 336 L= 21 W= 16